OpenStudy (anonymous):

in Ax+By=C... what do A, B, and C represent and how do you find those numbers?

OpenStudy (anonymous):

y= mx + b is the equation of the line with m as the slope b as the y-intercept. m and b can be derived by m=-A/B and b =C/B